OpenStudy (anonymous):

what is the derivative of this function ? 2x*sqrt(4-x)

zepdrix (zepdrix):

Start with product rule: \[\Large\sf \left(2x \sqrt{4-x}\right)'\quad=\quad \color{royalblue}{(2x)'}\sqrt{4-x}+2x\color{royalblue}{(\sqrt{4-x})'}\]There is the setup, need to take the derivative of the blue parts.

zepdrix (zepdrix):

This is a good derivative to memorize, it shows up a lot,\[\Large\sf \frac{d}{dx}\sqrt x=\frac{1}{2\sqrt x}\]

OpenStudy (anonymous):

i got stuck after taking the lcd

zepdrix (zepdrix):

Oh you wanted to combine the fractions to simplify?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\sf \left(2x \sqrt{4-x}\right)'\quad=\quad \color{royalblue}{(2x)'}\sqrt{4-x}+2x\color{royalblue}{(\sqrt{4-x})'}\]\[\Large\sf \left(2x \sqrt{4-x}\right)'\quad=\quad \color{orangered}{(2)}\sqrt{4-x}+2x\color{royalblue}{(\sqrt{4-x})'}\]\[\Large\sf \left(2x \sqrt{4-x}\right)'\quad=\quad \color{orangered}{(2)}\sqrt{4-x}+2x\cdot\color{orangered}{\frac{-1}{2\sqrt{4-x}}}\]Did you remember the negative that shows up from the chain rule?

OpenStudy (anonymous):

yes i did

zepdrix (zepdrix):

Cancel the twos and we're left to combine these two terms,\[\Large\sf \sqrt{4-x}-\frac{x}{\sqrt{4-x}}\]So it looks like the first term is missing a sqrt(4-x) in the denominator, yes? :o

zepdrix (zepdrix):

Ah I missed a two on the first term :d my bad

OpenStudy (anonymous):

once you put 2 on the first term is there anyway to simplify that more?

OpenStudy (anonymous):

\[\frac{ 2\sqrt{4-x} }{ \sqrt{4-x} }-\frac{ x }{ \sqrt{4-x}}\]

zepdrix (zepdrix):

First term needs a sqrt(4-x) for common denominator. So multiply the first term by sqrt(4-x) top and bottom:\[\Large\sf \color{green}{\left(\frac{\sqrt{4-x}}{\sqrt{4-x}}\right)}\cdot2\sqrt{4-x}-\frac{x}{\sqrt{4-x}}\]Yah you have the right idea, but you forgot to give the top a sqrt(4-x) also.

OpenStudy (anonymous):

do the sqrt cancel on the first term ?

OpenStudy (anonymous):

on the numerator i mean

zepdrix (zepdrix):

\[\Large\sf \color{green}{\left(\frac{\sqrt{4-x}}{\sqrt{4-x}}\right)}\cdot2\sqrt{4-x}-\frac{x}{\sqrt{4-x}}\]\[\Large\sf =\frac{2\sqrt{4-x}^2}{\sqrt{4-x}}-\frac{x}{\sqrt{4-x}}\]Cancel? Ehhh I guess if you want to think of it that way. You have a sqrt being squared away

zepdrix (zepdrix):

\[\Large\sf \frac{2(4-x)}{\sqrt{4-x}}-\frac{x}{\sqrt{4-x}}\]

OpenStudy (anonymous):

to get the critical number, i have to set this equal to zero right?

zepdrix (zepdrix):

Ya after you combine the fractions, \[\Large\sf 0=\frac{8-3x}{\sqrt{4-x}}\]Setting equal to zero will give us critical numbers.

OpenStudy (anonymous):

so x = 8/3?

OpenStudy (anonymous):

@zepdrix = \(\LaTeX\) master!

zepdrix (zepdrix):

Ok good. I guess we would also want to include values which are undefined in the derivative. sqrt(4-x) in the denominator gives us another zero. it's a different kind of critical point, but uhh we should probably still include it.

zepdrix (zepdrix):

lol :3

zepdrix (zepdrix):

Setting the denominator equal to zero, \[\Large\sf \sqrt{4-x}=0\]What value of x do we get here? :o

OpenStudy (anonymous):

4

OpenStudy (anonymous):

?

zepdrix (zepdrix):

yes :)

OpenStudy (anonymous):

would that also be a critical point ?

OpenStudy (anonymous):

and also, if you graph this... how would you find that absolute max?

zepdrix (zepdrix):

Yes. It's not the same type of critical point as x=8/3. It's telling us where the function has a `vertical` tangent which is not as helpful. But it's still a value we would consider when trying to determine max/min and such.

zepdrix (zepdrix):

I guess we would want to determine what type of critical point x=8/3 is. :o

zepdrix (zepdrix):

There are several ways to do that. I like doing this personally. Drawing out a number line and plugging in a couple of test points to see what's happening on the left and right of our critical number.

OpenStudy (anonymous):

right, which ive already solved. so to get that critical point which is also your abs and local max, you ust plug 8/3 to the equation ?

zepdrix (zepdrix):

Oh so you can write it as an ordered pair? Yah that sounds good! Back into the `original equation`.

OpenStudy (anonymous):

ok cool . thanks so much!